2 attempts left Check my work Convert 1595 mm Hg to Pa. x 10Pa
Answer #1

we know 1atm=760mm of Hg

thus 1595mm of Hg= 1595/760= 2.098atm

now 1atm= 1.013 * 10 Pa

2.098atm=2.098 * 1.013 * 10Pa=2.125 * 10°Ρα